FEL Upgrade Project Weekly Brief

March 4-8, 2002

 

 

Highlights:

Completion of the gun chamber and leak check of the ceramic stack highlighted significant steps forward for the injector gun this week.

The optic systems made major headway with the release of drawings for the alignment assembly and signoff of a requisition for a ultrafast diagnostic laser.

And last but not least in the transport area we potted a complete coil pair for the optical chicane dipoles and all 3" quadrupoles for the IR system are now in hand and meet spec.

Management:

With the completion of several minor adjustments to the baseline budget of the IR Upgrade program and the UV upgrade program, we completed and distributed the project monthly reports for December 2001 and January 2002. The baseline budget changes in the IR upgrade project included requested scope changes recommended by the project review committee and new total project budget (from $4,500k to $4,442k for Phase 2) that was adjusted to reflect the actual budget available for the Phase 2 project. The total project budget for the Phase 1 UV project is $2,833k. With the finalization of the largest subcontract in this project (to the Aerospace Corp.), we could fix the remaining budget lines for the Jefferson Lab activities. We expect subsequent monthly reports to return to their usual schedule of 15-20 days after the close of the month.

On Wednesday March 7 we made a requested visit with the RDT&E subcommittee staff of the Senate Armed Services Committee , which was a follow-up to visits from this committee to the Jefferson Lab for a briefing on the FEL program in December and January. We learned that a visit is tentatively scheduled by members of the RDT&E subcommittee staff of the House Armed Services Committee on next Friday (March 15th).

On Friday March 8, we were visited by the Dr. Reinhart Poprawe, Managing Director of the Fraunhofer Institute of Laser Technology (ILT) in Aachen, Germany. The ILT has a unique international reputation in laser materials processing. Their program is 50 % supported by industry. We discussed several areas of potential collaboration.

WBS 3 (Beam Physics):

Analysis of GX, GQ and GY field quality underwent a femto-review; see WBS 9 - Biallas for results and actions.

A simple geometric description of the UV trajectory through the GX was derived (deranged?).

Documentation of endloop dipole considerations was started.

WBS 4 (Injector):

The gun chamber has been completed. The ceramic stack was leak checked ok with N.I.L. The cathode support tube was mounted into W&M's implantation system and will take about two weeks to complete. An engineering review of the resistor was held, minor changes are being implemented.

Gun HVPS - A design review for the combined running and conditioning resistors was held this week. Agreement was finally reached on how to join the 2 major pieces. The existing probes and fittings will be used to operate the internal motor and check the resistance.

DC Field Emission scan of a previously hydrogen cleaned wafer was completed at 60 MV/m. Three emitters were found at 54, 47 and 30 MV/m of applied electric field. The maximum observed current was around 2 nA. The current was observed to be very erratic while varying the applied voltage to obtain the field enhancement factor, which was found to be between 30 and 60. The data are being analyzed to estimate the effect of Cesium on the emission current. An EDX spectrum was taken and Indium was found on one of the emitters. The rest of the wafer shows only Gallium and Arsenic as expected.

WBS 6 (RF):

Quarter HVPS - Low level wiring continues. System prints are being updated with the work. The 5 uf 40 kV capacitors were received this week.

Quarter Klystrons - The drawings for the klystron carts were completed and signed this week. Two additional klystron carts will be made by our machine shop. The first 12' of main LCW piping in the RF Gallery is being replaced to increase flow and to add the larger taps for the upgrade. This effort will take several weeks to complete.

WBS 8 (Instrumentation):

The testing apparatus for the beamviewers and cameras is being designed with testing to start soon.

Progress continues on the Beam Loss Monitoring System upgrade, the number and locations of monitors has been defined. Designs were completed on the Breakout Module and turned over to EECAD for documentation.

A VME digital interface card has been designed and prototypes are in the procurement phase.

WBS 9 (Transport):

Dipoles

Optical Chicane Dipoles (GW)

At Magnet Enterprises International in Oakland CA. the second coil was potted successfully, making a complete pair for the first article magnet.

Return legs are being ground to their final size for a projected finish on Monday. Top and bottom slabs are finished in heat treat.

Field clamps are in process. Covers are in. Manifolds are being shipped.

Injector Dipoles (DU/DV)

Master Machine picked up material and is torch cutting it into the pieces for manufacturing.

We signed off the revision A drawings and have distributed them to both Master Machine (Cores) and Wang NMR (Coils) for incorporation.

Arc Dipoles (GY, GX, GQ)

Electronic procurement of Bend (GX) and Reverse Bend (GQ) Dipoles continued.

David Douglas completed his analysis of the use of the Bend and Reverse Bend Dipoles (GX & GQ) as well as the 180 Dipole (GY) at all levels of excitation using the magnetic model analysis from AES. The conclusion is that the magnets meet specification at 210 MeV/c but have to be modified to be used successfully throughout the full range of excitation anticipated (a lower limit of 80 MeV/c). Modifications are:

- Add separate, ganged 2500 G-cm correction coils to the GQ dipole coils astride the existing correction coil for use in matching the dipoles to the GX magnets at low energy.

- Shave the length of the GQ by 5 mm and the GX by 3 mm to bring the field integrals down to match the core field at low energy. We will expand the length to the high energy requirement using shims.

- The UV operation of the GX may need both plus and minus ability as well as shortened magnet. Through that arc we will maintain the 3 mm decrease in length and place a trim power supply on those portions of the coils that are switched out during UV operation.

Wang NMR of CA continues making the GY coil winding mandrel and they are

designing the potting mold for the GY coils.

Bosma Machine has the GY core pieces in heat treatment.

Quadrupoles

3 inch quad (QX)

All QX quadrupoles needed for the IR portion of the FEL Upgrade are now on hand.

Production measurements were taken on 25 QX Quadrupoles with no evident rejections.

We are looking to release a batch of magnets next week after review and sign off by David Douglas.

Trim Quad (QT)

Milhous Control of Virginia is making the side coils with the material we recently

supplied them.

The wire manufacturer is taking back the material for the top and bottom coils for re spooling without a twist. A twist in the original reels of copper made it impossible to wind successfully.

Sextupole (SF)

DULY Research has started analysis again, lengthening the magnet to bring the saturation in the pole tips down.

Octupole (OT)

Work on this magnet is on furlough until the sextupole is designed.

Beam Line and Vacuum

AES continues drawing up the chambers throughout the arcs.

Design of the Arc Dipole Stands continued.

Design continued on the details for the region after the cryomodules.

Butch Dillon-Townes finished updating the schedule and passed it on the scheduling folks for incorporation into the master schedule.

WBS 10 (Wiggler):

Measurements of the dispersion section show good field quality but the corrector needs to be 50% higher than previous measurements. We are trying to track down why this is so. The field integral without the corrector is a factor of two larger than previously.

WBS 11 (Optics):

The drawings for the mirror metrology system (OCMMS) were signed and released for procurement. Work began on control of the piezos that will actually do the fast steering of the optical cavity mirrors. This week a LabView program was written that can read the signal from the position sensitive detector, and output a control signal to the piezo amplifier. The amplifiers will arrive next week and we will then work to control them.

We continue work on the insertable mirror drawings. We are working on the vacuum vessel drawings. We held meetings on limit switches for the optical cavity assemblies and nomenclature for the entire optical transport system. As a result of the latter meeting, updated drawings are being created. A spreadsheet detailing the vacuum wiring for the various optical assemblies (optical cavity, collimator and turning mirror cassettes) was generated and given to I&C so they can have electrical schematics produced and order components such as vacuum feedthroughs. We began collecting more data on the deformable mirror assembly in the mirror test stand (MTS), with the goal of studying the ROC as a function of water flow rate. We also spent some time determining the sensitivity of the ROC measurement, using the laser interferometer, to software and hardware parameters.

A requisition for an ultrafast laser system was signed this week. This laser will be used for FEL output and electron beam physics studies.

Progress was also made in compressing the pulsewidth of the "waste" drive laser IR output, using nonlinear effects in a single-mode fiber. So far, we have determined that the damage threshold is greater than 70 MW/cm^2, in excess of the manufacturer's claim of 50 MW/cm^2.

UV FEL Progress

Review of the last 10 years literature on laser induced damage thresholds in the UV and DUV is underway.

 

 

--Fred/dmag